OpenStudy (anonymous):

If the edges of a cube are increased by 2 inches, A) the volume is increased by 8 cubic inches B) the area of each face is increased by 4 square inches C) the diagonal of each face is increased by 2 inches D) the sum of the edges is increased by 24 inches

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

there're 12 edges in a cube. let edge be, x then sum of edges (let be, A)= 12x given, if edge is increased by 2, i.e. x becomes (x+2) then new sum of edges= 12(x+2) = 12x + 24 = sum of edges (A) + increment clearly, option D. is correct.

OpenStudy (anonymous):

ooohhhh i see it now. thanks!

OpenStudy (anonymous):

cheers!